# gather+[1][1]

### v1

@@ 6       $\mathcal{T}_{in}\leftarrow\mathcal{T}_{(d)}$
8       for each inter-device partition, $iDp_{d,z}$ in $\mathcal{T}_{in}$, $z=0,1\ldots,k_{d}-1$ parallel do
9             // Execute threadblocks in parallel across GPU SMs
10             for each inter-SM partition $iSMp_{d,z,b}$ in $iDp_{d,z}$, $b=0,1,\ldots,t_{d,z-1}$ parallel do
11                   $nnz\leftarrow 0$
12                   for $nnz<|iSMp_{d,z,b}|$ parallel do
13                         for each column, $l$ in threadblock parallel do
14                               if $nnz+l<|iSMp_{d,z,b}|$ then
15                                     Load($x_{i}$ at $(nnz+l)$)
16                                     $\hat{F}_{d,z}\leftarrow$OFMRU($x_{i}$,F)
17
18
19                        // $\theta$ = number of columns in a threadblock
20                         $nnz\leftarrow nnz+\theta$
@@
23            __Inter-GPU Barrier__
24             // Update output factor matrix copy on each GPUs
25             // See Algorithm 3 on GPU to GPU communication
26             $F_{d}\leftarrow\hat{F}_{d,z}$
27
28      __Inter-GPU Barrier__
@@ 26             $F_{d}\leftarrow\hat{F}_{d,z}$
Algorithm 1 Overall proposed algorithm on multi-GPUs for a single iteration
Algorithm 1 shows the overall parallel algorithm to execute MTTKRP across all modes of a single iteration. Algorithm 1 takes (1) all copies of the $N$ mode input tensor $\textbf{T}=\{\mathcal{T}_{(0)},\mathcal{T}_{(1)}\cdots\mathcal{T}_{(N-1)}\}$ and (2) the factor matrices $\textbf{F}=\{F_{0},F_{1},...,F_{N-1}\}$ computed in the previous iteration. MTTKRP is performed mode by mode (Algorithm 1: line 4). Each GPU operates on the inter-device partitions ($iDps$) statically assigned by the proposed load balancing scheme in Section 4.9 prior to computation.
Once an $iDp$ is assigned to a GPU for computation, the GPU operates on $iDp$ independently of other GPUs (Algorithm 1: line 8). Note that $iDps$ do not have task dependencies between them and can be executed independently (see Observation 1). Each GPU SM of the GPU operates on an inter-SM partition ($iSMp$) in the assigned $iDp$ (Algorithm 1: line 9-10).
In Algorithm 1, $iSMp_{d,z,b}$ denotes the $b^{\text{th}}$ inter-SM tensor partition of the $z^{\text{th}}$ $iDp$ in mode $d$ (see Section 3.2). When a GPU SM is available for computation, a threadblock and its associated inter-SM tensor partition are assigned to the GPU SM. The threadblock performs OFMRU on each nonzero tensor element in the assigned $iSMp$ (Algorithm 1: line 18). To support OFMRU, we use a threadblock with dimensions of $R\times\theta$ executed on each GPU SM. Here, $R$ denotes the rank of factor matrices, and $\theta$ indicates the number of nonzero tensor elements loaded to a threadblock at once. $R\times\theta$ is kept equal to the number of threads in a GPU SM to utilize all the threads in a GPU SM. Each column in the threadblock loads a nonzero tensor element at a time and distributes it across the threads in the same column. Each thread in the column extracts the embedded information from the loaded nonzero tensor element (Algorithm 1: lines 12-18). and performs OFMRU (Algorithm 1: line 16).
After processing $iDp$, the updated factor matrix rows are shared across the GPUs (Algorithm 1: line 22) and the local copy of each GPU is updated before executing MTTKRP for the next output mode.
1 OFMRU($x_{i}$, F):
2 Input: Nonzero tensor element $x_{i}$
3           Factor matrices $\textbf{F}=\{F_{0},F_{1},...,F_{N-1}\}$
4 Output: Updated $F_{d,z}$
5 $value\leftarrow val_{i}$
6 $\beta_{i}=(c_{0},\ldots,c_{N-1})$
7
8// $\ell$ is a vector of size R
9 for each rank $r=0,\ldots,R-1$ parallel do
10       $\ell(r)\leftarrow value$
11
12for input mode $w\in\{0,\ldots,N-1\}\setminus\{d\}$ do
13
14      $vec\leftarrow$ Load(row $c_{w}$ from $w^{\text{th}}$ factor matrix, $F_{w}$)
15       // Row 00 to $R-1$ of the threadblock perform independent computations
16       for each rank $r=0,\ldots,R-1$ parallel do
17             $\ell(r)\leftarrow\ell(r)\times vec(r)$
18
19
20
21for each rank $r=0,\ldots,R-1$ parallel do
22       $Y_{d}(c_{d},r)\leftarrow\text{Atomic}(Y_{d}(c_{d},r)+\ell(r))$
23
Algorithm 2 Update the factor matrix row of mode $d$
4.6. Parallel Algorithm of Output Factor Matrix Row Update Computation
Algorithm 2 describes the Output Factor Matrix Row Update Computation (OFMRU) on each nonzero tensor element. In Algorithm 2, the rows of the input factor matrices are loaded from the global memory of the GPU (Algorithm 2: line 11) depending on the indices of the current nonzero tensor element ($x_{i}$) in the GPU thread. Subsequently, the SM threads perform OFMRU (see Section 3.1) using the nonzero tensor element and the rows of the factor matrices (Algorithm 2: lines 10 - 16). To enable threadwise parallelism in OFMRU, each thread in a column performs computations on its respective rank (Algorithm 2: lines 12 - 14). Each GPU threadblock updates the output factor matrix (Algorithm 2: lines 16), using atomic computations to avoid race conditions between threadblocks in the same GPU.
@@ 4.7. Host CPU-GPU Communication
The input tensor copies are initially stored in the host CPU memory. The inter-device partitions (iDps) are transferred from the host CPU memory to each GPU global memory during the execution time of each output mode.
4.8. GPU to GPU Communication
1 P2P Communication ($gpu\_id$, $F_{d,gpu_{i}d}$):
@@
5       Send_Copy = $F_{d,(gpu\_id+z)modM}$
6       Send(Send_Copy, $(gpu\_id+1)modM$, size(Send_Copy))
7       Rec_Copy = $F_{d,(gpu\_id-z-1)modM}$
8       Receive(Rec_Copy, $(gpu\_id-1)modM$, size(Rec_Copy)))
@@ 9       Barrier
For a given output mode, once a GPU completes processing the iDp assigned to it, the GPU communicates the updated rows to all the GPUs in the platform. The distributed factor matrix will act as an input factor matrix in the following modes of MTTKRP computations.
We adopt the ring network communication model (Li et al., 2018b) to broadcast the partitions of the factor matrices among the GPUs, as shown in Algorithm 3. We use GPU to GPU communication to send and receive partitions of the factor matrices across GPU without involving the CPU host memory. The ring network model is suitable for bulk transfers among neighboring devices with limited bandwidth, which is ideal for communicating the output factor matrix among multiple GPUs.
4.9. Load Balancing
After assigning an iDp to a GPU, the number of OFMRU operations performed by the GPU on the iDp is proportional to the number of nonzero tensor elements in the iDp. The number of nonzero tensor elements in each iDp varies according to the sparsity of the tensor. In our work, the workload is distributed by assigning iDps of different sizes to GPUs, aiming to minimize the difference between the number of nonzero elements assigned to each GPU for computation.
Consider an input tensor divided into iDps $\{\textit{iDp}_{d,j}:\forall\,j\}$ for an output mode $d$. Let $\eta_{d,j}$ denote the number of nonzero tensor elements in $\textit{iDp}_{d,j}$. We arrange iDps so that $\eta_{d,j}\geq\eta_{d,j^{\prime}}$ if $j<j^{\prime}$, thus ordering the iDps in descending order according to the number of nonzero tensor elements. Each iDp is iteratively assigned to the GPU with the fewest number of nonzero tensor elements assigned at the time. This procedure is repeated for all modes of the tensor.
For a given output mode, let $\eta_{\textrm{max}}$ be the largest number of nonzero tensor elements assigned to a single GPU and $\eta^{*}_{\textrm{max}}$ be the maximum number of nonzero tensor elements assigned to a GPU in the optimal distribution possible for the given iDps. Then our proposed solution guarantees $\eta_{\textrm{max}}\leq 4/3\cdot\eta^{*}_{\textrm{max}}$ (Wijeratne et al., 2023b; Graham, 1969).
@@ 5.2. Overall Performance
For evaluation, we set the rank of the factor matrices ($R$) to 32, similar to the state-of-the-art (Nisa et al., 2019c; Nguyen et al., 2022a; Li et al., 2018a). Our work demonstrates a geometric mean speedup of 5.1$\times$ compared with the state-of-the-art baselines.
When out-of-memory computation is enabled, BLCO (Nguyen et al., 2022a) stores the input tensor in the host CPU memory and loads the tensor partition by partition from the external memory of the host CPU to the GPU global memory. In BLCO, using a single GPU introduces additional memory traffic between the host CPU and the GPU. Meanwhile, our work has more effective bandwidth between the host CPU and the GPU since multiple GPUs can concurrently communicate with the host CPU. Our work shows a geometric speedup of 5.1×5.1$\times$5.1 × using 4 GPUs compared to BLCO.
@@
FLYCOO-GPU (Wijeratne et al., 2024) does not support Amazon, Patents, and Reddit since these tensors do not fit in the GPU global memory. FLYCOO-GPU requires maintaining 2 copies of the tensor in the GPU global memory. On Twitch, FLYCOO-GPU outperforms our work by 3.9$\times$ due to the communication overhead of our work. Twitch is the smallest billion-scale tensor in the literature. The small size enables keeping 2 tensor copies inside the global memory of a single GPU. Since FLYCOO-GPU only targets a single GPU, FLYCOO-GPU does not require GPU-GPU communication or host CPU-GPU communication.
Figure 4. Sustained compute performance on different number of GPUs
5.3. Sustained Compute Performance
Figure 4 shows the sustained compute performance of our work on different numbers of GPUs. In all tensors, we observe a nearly linear increase in sustained compute performance since the computation time (i.e., each GPU SM performing OFMRU and GPU global memory accesses) is the dominating factor in execution (see Section 5.5).
5.4. Impact of Load Balancing Scheme
Figure 5. Impact of proposed load balancing scheme
Section 4.9 describes the load balancing scheme used in this work to distribute the inter-device partitions among GPUs.
An alternative approach is to distribute the non-zero tensor elements equally among all GPUs. It introduces additional computations on the host CPU to merge the partial results of each inter-device partition. Figure 5 compares our proposed strategy with the equal distribution of nonzero tensor elements among GPUs. Our proposed load balancing scheme achieves 5.3$\times$ to 10.3$\times$ speedups in total execution time.
@@ 5.5. Execution Time Breakdown
For large tensors such as Patents and Reddit, moving inter-device partitions from CPU host memory to GPU global memory is the major contributor to the total communication time. Tensors with a large number of indices (e.g., Amazon and Twitch) require frequent GPU-GPU communication to update factor matrices at the end of each output mode computation, which significantly contributes to the total communication time. In particular, Reddit exhibits a significant communication overhead ($32\%$) during the execution due to (1) the large number of nonzero tensor elements in the tensor and (2) the modes with a large number of indices.
5.6. Load Balancing Workload among GPUs
@@
As discussed in Section 4, each GPU updates a set of rows of the output factor matrix. In this section, the computation time of a GPU is defined as the total time to perform OFMRU on all nonzero tensor elements assigned to a GPU across all modes. The computation time overhead is the difference between the maximum and minimum computation times among GPUs on the target platform.
Figure 7 shows the computation time overhead among GPUs (as a percentage) on our target 4 GPU platform. To determine the computation time of each GPU, we execute each GPU grid (see Section 4.2) separately and measure the computation time in each output mode. The computation time overhead for all billion-scale tensors is less than $1\%$ (as a percentage of the total time required to perform all OFMRU computations using all 4 GPUs).
@@ 5.7. Scalability
Figure 8 shows the speedup of each input tensor as the number of GPUs increases from 1 to 4. The speedup is calculated by dividing the total execution time of the input tensor on a single GPU by its total execution time in each case. Our proposed parallel algorithm achieves geometric mean speedups of 1.9$\times$, 2.3$\times$, and 3.3$\times$ when using 2, 3, and 4 GPUs, respectively, compared to a total execution time of a single GPU. The speedup increases nearly linearly with the number of GPUs for each tensor.
@@ 6. Conclusion and Future Work
In this paper, we proposed AMPED, a novel parallel algorithm designed to accelerate MTTKRP on billion-scale tensors using multiple GPUs. Our approach introduced a partitioning scheme that allowed each tensor partition to be executed independently of other partitions in each output mode. The partitioning scheme, coupled with the proposed load-balancing strategy to distribute the workload across all the GPUs, minimized the GPU idle time. AMPED achieved a geometric mean speedup of 5.1$\times$ (using 4 GPUs) in total execution time compared with the state-of-the-art GPU baselines.

### v2

@@ 5 $T_{in}\leftarrow T_{d}$
7 for each tensor shard, $TS_{d,z}$ in $T_{in}$ parallel do
8    $\hat{Y}_{d}$ $\leftarrow$ GPU Grid($TS_{d,z}$, Y)
9
10 __Inter-GPU Barrier__
11 // Exchange generated output factor matrix partition across GPU
12 $Y_{d}$ $\leftarrow$ All Gather($\hat{Y}_{d}$)
13 __Inter-GPU Barrier__
14
Algorithm 1 Overall Proposed Algorithm
After processing TSTS, the updated factor matrix rows are shared across the GPUs (Algorithm 1: line 10) using all-gather communication primitive (Kobus et al., 2019) before executing MTTKRP for the next output mode.
4.7. Mapping Parallel Algorithm to GPU threadblocks
Figure 4. Overview of threadblock
1 GPU Grid($TS_{d,z}$, Y, $T_{out}$):
2 Input: Input tensor shard, $TS_{d,z}$
3 Factor matrices $\textbf{Y}=\{Y_{0},Y_{1},...,Y_{N-1}\}$
4
5Output: Updated factor matrix of mode $d$, $\hat{Y}_{d}$
6
7// Execute threadblocks in parallel across GPU SMs
8 for each inter-shard partition $ISP_{d,z,b}$ in $TS_{d,z}$, $b=0,1,\ldots,t_{d,z-1}$ parallel do
9 for each column, $t$ in threadblock parallel do
10    if $nnz+t<|ISP_{d,z}|$ then
11       Load($x_{i}$ at $(nnz+t)$)
12       $value\leftarrow val_{i}$
13       $p_{i}=(c_{0},\ldots,c_{N-1})$
14       // Elementwise Computation
15       for input mode $w\in\{0,\ldots,N-1\}\setminus\{d\}$ do
16
17          $vec\leftarrow$ Load(row $c_{w}$ from $w^{\text{th}}$ factor matrix)
18          // Row 0 to R1R-1 of the threadblock perform independent computations
19          for each rank $r$ in $R$ parallel do
20             $\ell(r)\leftarrow\ell(r)\times vec(r)$
@@
23
24    for each rank $r$ in $R$ parallel do
25       $\hat{Y}_{d}(c_{d},r)\leftarrow\text{Atomic}(\hat{Y}_{d}(c_{d},r)+\ell(r))$
26
27 // P is the number of columns in a threadblock
28 $nnz\leftarrow nnz+P$
@@
30
Algorithm 2 Parallel Algorithm Executed on a Shard
The basic computing unit of a GPU is a thread. According to the GPU programming model (Ansorge, 2022; Zeller, 2011), a multi-threaded program is partitioned into blocks of threads (i.e., threadblocks) that operate independently.
In our proposed algorithm, a threadblock has a dimension of $R\times P$, where $R$ denotes the rank of factor matrices and $P$ indicates the number of nonzero tensor elements parallelly loaded to a threadblock (see Figure 4). Here, each column of the threadblock shares the same nonzero tensor element, and each column performs elementwise computation on a nonzero tensor element.
Algorithm 2 outlines the computations executed on a shard partition. In Algorithm 2, $ISP_{d,z,p}$ corresponds to $p^{\text{th}}$ inter-shard partition in $z^{\text{th}}$ tensor partition of mode $d$. When a GPU SM is idle, a threadblock and its corresponding inter-shard partition are assigned to the GPU SM for computation. Each column in the threadblock loads a single nonzero tensor element at a time and shares the nonzero tensor element across the threads in the same column. Each thread in a column extracts the information from the tensor element $x_{i}$ in COO format (Algorithm 2: lines 9-11). Subsequently, each threadblock performs elementwise computation (Algorithm 2: lines 6-17). To achieve threadwise parallelism, each thread in a column only executes the update operation on a single column of a row of the output factor matrix (Algorithm 2: lines 15 - 17). The rows of the input factor matrices are loaded from the GPU global memory (Algorithm 2: lines 13-14) depending on the indices of the current tensor element ($p_{i}$) executed in the GPU thread. Each GPU threadblock locally updates the output factor matrix (Algorithm 2: lines 15) while maintaining the coherence of each threadblock to ensure the correctness of the program. According to the proposed partitioning scheme, threadblocks require atomic operations for tensor elements that reside in the same tensor shard. Hence, we use atomic operations across the GPU threadblocks within the same GPU to maintain the correctness of the program (Algorithm 2: lines 18-19).
@@ 4.8. Host CPU-GPU Communication
The input tensor copies are initially stored in the host CPU memory. The tensor shards are transferred from the host CPU memory to each GPU global memory during the execution time of each output mode.
4.9. All Gather Communication (GPU to GPU)
For a given output mode, once a GPU completes processing the tensor shard (TS) assigned to it, the GPU communicates the updated rows to all the GPUs in the platform. The distributed factor matrix will act as an input factor matrix in the following modes of MTTKRP computations.
We adopt the ring network communication model (Li et al., 2018b) to perform all-gather among the GPUs, as shown in Algorithm 3. We use GPU-to-GPU communication to send and receive partitions of the factor matrices across GPUs without involving the CPU host memory. The ring network model is suitable for bulk transfers among neighboring devices with limited bandwidth, which is ideal for communicating the output factor matrix among the GPUs.
1 All Gather ($\hat{Y}_{d,gpu\_id}$):
2 Input: Mode $d$ output factor matrix partition in the local GPU, $\hat{Y}_{d,gpu\_id}$
@@
7   Send_Copy = $\hat{Y}_{d,(gpu\_id+z)modM}$
8   Send(Send_Copy, (gpu_id+1)modM(gpu\_id+1)modM, size(Send_Copy))
9   Rec_Copy = $\hat{Y}_{d,(gpu\_id-z-1)modM}$
10   Receive(Rec_Copy, (gpu_id1)modM(gpu\_id-1)modM, size(Rec_Copy)))
@@ 5.2. Overall Performance
[…]
When out-of-memory computation is enabled, BLCO (Nguyen et al., 2022a) stores the input tensor in the host CPU memory and loads the tensor partition by partition from the external memory of the host CPU to the GPU global memory. In BLCO, using a single GPU introduces additional memory traffic between the host CPU and the GPU. Meanwhile, our work has more effective bandwidth between the host CPU and the GPU since multiple GPUs can concurrently communicate with the host CPU. Our work shows a geometric speedup of $5.1\times$ using 4 GPUs compared to BLCO.
@@
[…]
5.3. Impact of Partitioning Scheme
Figure 6. Impact of proposed partitioning scheme
Section 3 describes the partitioning scheme used in this work to distribute the tensor shards among GPUs.
[…]
@@ 5.4. Execution Time Breakdown
For large tensors such as Patents and Reddit, moving tensor shards from CPU host memory to GPU global memory is the major contributor to the total communication time. Tensors with a large number of indices (e.g., Amazon and Twitch) require frequent GPU-GPU communication to update factor matrices at the end of each output mode computation, which significantly contributes to the total communication time. In particular, Reddit exhibits a significant communication overhead ($32\%$) during the execution due to (1) the large number of nonzero tensor elements in the tensor and (2) the modes with a large number of indices.
@@
5.5. Workload Distribution among GPUs
@@
As discussed in Section 4, each GPU updates a set of rows of the output factor matrix. In this section, the computation time of a GPU is defined as the total time to perform elementwise computation (EC) on all nonzero tensor elements assigned to a GPU across all modes. The computation time overhead is the difference between the maximum and minimum computation times among GPUs on the target platform.
Figure 8 shows the computation time overhead among GPUs (as a percentage) on our target 4 GPU platform. To determine the computation time of each GPU, we execute each GPU grid (see Section 4.2) separately and measure the computation time in each output mode. The computation time overhead for all billion-scale tensors is less than $1\%$ (as a percentage of the total time required to perform all EC using all 4 GPUs).
@@ 5.6. Scalability
[…]
@@ 6. Conclusion and Future Work
[…]